A contractor finds it takes Sam 11hours to construct a wall of certain size. It takes Jime 6hours to do it. How long would it take if the work together?
Answer by Alan3354(69443)

A contractor finds it takes Sam 11hours to construct a wall of certain size. It takes Jime 6hours to do it. How long would it take if the work together?
-------------------
The "trick" to this type of problem is to find terms that can be added.
Adding 11 and 6 is obviously not it.
Sam takes 11 hrs to do a wall, so in 1 hour he does 1/11 of a wall.
Jime does 1/6 of a wall per hour.
Add 1/6 and 1/11 = 17/66 together.
If they do 17/66 of a wall per hour, it will take 66/17 hours to do a wall.
66/17 hours is it.
---------------
There's a shortcut you can use, product/sum, for 2 workers.
11*6/(11+6) = 66/17